OpenStudy (anonymous):

if k is a positive integer between 0 and 60 what is one possible value of k for which Square root(k+7)-5 is a posstive integer?

OpenStudy (anonymous):

\[\sqrt{(k+7)-5}\]

OpenStudy (anonymous):

Have the psat's tomorrow so need help with this question.

OpenStudy (anonymous):

Is the 5 under the square root or not? You wrote it differently.

OpenStudy (anonymous):

Assuming it isn't, there are a few ways to approach the problem. The fastest would probably be to see that you want sqrt(k+7) > 5 and an integer. So, pick 6 (since you want k fairly small if it's between 0 and 60). So, 6 = sqrt(36) = sqrt(k+7) So, 36 = k + 7. => k = 29. And, you're done.

OpenStudy (anonymous):

its under the square root